# Block-Diagonal Orthogonal Relation and Matrix Entity for Knowledge Graph Embedding

### v1

@@ Abstract
The primary aim of Knowledge Graph embeddings (KGE) is to learn low-dimensional representations of entities and relations for predicting missing facts. While rotation-based methods like RotatE Sun et al. (2019) and QuatE Zhang et al. (2019) perform well in KGE, they face two challenges: limited model flexibility requiring proportional increases in relation size with entity dimension, and difficulties in generalizing the model for higher-dimensional rotations. To address these issues, we introduce OrthogonalE, a novel KGE model employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. This approach enhances the generality and flexibility of KGE models. The experimental results indicate that our new KGE model, OrthogonalE, is both general and flexible, significantly outperforming state-of-the-art KGE models while substantially reducing the number of relation parameters. Our code is available at OrthogonalE
[…]
@@ 1 Introduction
The fundamental elements of knowledge graphs (KGs) are factual triples, each represented as $(h,r,t)$, indicating a relationship $r$ between head entity $h$ and tail entity $t$. Notable examples include Freebase Bollacker et al. (2008), Yago Suchanek et al. (2007), and WordNet Miller (1995). KGs have practical applications in diverse fields such as question-answering Hao et al. (2017), information retrieval Xiong et al. (2017), recommender systems Zhang et al. (2016), and natural language processing Yang and Mitchell (2019), garnering considerable interest in academic and commercial research.
@@
Recently, rotation-based KGE methods have achieved notable success in the field. For instance, RotatE Sun et al. (2019) conceptualizes relations as 2D rotations, while QuatE Zhang et al. (2019) employs 3D rotations. Essentially, as illustrated in Fig. 1 , both operate by multiplying the relation matrix $\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}$ composed of numerous small rotation blocks $\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$ (RotatE: $\mathbb{R}^{2\times 2}$, QuatE: $\mathbb{R}^{3\times 3}$) with the head entity vector $\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}$.
Figure 1: Fundamental operations and inherent challenges of rotation-based KGE models.
@@
Nevertheless, these approaches encounter two primary issues as shown in Fig. 1. First, the model’s lack of flexibility necessitates increasing the overall relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{(n+l)\times(n+l)}$) to meet entity dimension requirements ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{n+l}$) for better represent entities, which often leads to redundancy and inefficiency in representing relations. Second, exploring high-dimensional rotational KGE models is challenging due to the significant computational demands and complexity of rotations in higher dimensions ($\mathbf{B}_{i}:\mathbb{R}^{2\times 2},\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{d\times d},d>3$), such as 4D and 5D rotation. This restricts the development of more generalized and higher-dimensional rotation KGE approaches.
@@
To overcome these issues, we propose a highly general and flexible KGE model named OrthogonalE as shown in Fig. 3. Firstly, by transforming entity vectors $\mathbf{e}_{h}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{H}\in\mathbb{R}^{n\times m}$, we control the entity dimension through variable $m$, avoiding unnecessary expansion of the relation size. Secondly, leveraging the concept that rotation matrices are orthogonal, we replace rotation matrices $\mathbf{B}_{i}$ with orthogonal matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ of adaptable dimensions $d$, facilitating the exploration of higher-dimensional block-diagonal orthogonal matrix models. Lastly, for effective optimization, we employ Riemannian optimization for the relation matrix $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ and Stochastic Gradient Descent (SGD) for the entity matrix $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$.
@@
We evaluate the new model on two KGE datasets including WN18RR Dettmers et al. (2018), FB15K-237 Toutanova and Chen (2015). The experimental results demonstrate that our new KGE model, OrthogonalE, is both general and flexible, significantly outperforming state-of-the-art KGE models while substantially reducing the number of relation parameters.
@@ 2 Related Work
Translation-based approaches are prominent in KGE, notably TransE Bordes et al. (2013), which interprets relations as vector translations. TransH Wang et al. (2014), TransR Lin et al. (2015), and TransD Ji et al. (2015) represent extensions of the translation-based method, building upon the foundational approach of TransE. ComplExTrouillon et al. (2016) advances this by embedding entities and relations in a complex space and using the Hermitian product for modeling antisymmetric patterns, a technique that has proven highly effective in learning KG representations. Inspired by ComplEx, RotatE Sun et al. (2019) then innovated by treating relations as rotations in a complex vector space, capable of capturing varied relation patterns like symmetry and inversion. Following this, QuatE Zhang et al. (2019) employed quaternion operations (3D rotations) for even better expressiveness than RotatE.
However, considering the two major disadvantages of rotation-based methods mentioned in the Introduction 1, there is a need to refine our model to make it more general and flexible.
@@
Riemannian Optimization has concurrently advanced in deep learning domains, particularly in CNNs and RNNs. Cho and Lee (2017) innovatively substituted CNN’s Batch Normalization layers with Riemannian optimization on the Grassmann manifold for parameter normalization. Additionally, significant strides in stabilizing RNN training have been made by Vorontsov et al. (2017), Wisdom et al. (2016), Lezcano-Casado and Martınez-Rubio (2019), and Helfrich et al. (2018), through the application of Riemannian optimization to unitary matrices.
@@ 3.2 Optimization on the orthogonal manifold
To address problem 1 more efficiently, recent studies suggest optimization on the orthogonal manifold with retractions as an effective approach Ablin and Peyré (2022). In this work, we primarily employ the retraction with exponential map method for iterative optimization, as illustrated in Figure 2. The key iteration formula for this method is:
@@
Where $\text{expm}()$ denotes the matrix exponential.
@@ 4 Approach
Figure 3: Diagram of the OrthogonalE approach.
Our approach aims to acquire a flexible and general KGE model with employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations.
@@ 4.1 Orthogonal Matrices for Relations
RotatE (2D rotation) and QuatE (3D rotation) inherently employ fixed rotation matrices ($\mathbf{B}_{i}\in\mathbb{R}^{d\times d},d\leqslant 3$), as depicted in Figure 1. However, the exploration of high-dimensional rotational KGE models ($\mathbf{B}_{i}\in\mathbb{R}^{d\times d},d>3$) is limited by substantial computational demands and the complexity of rotations in higher dimensions, which hinders the advancement of more generalized, high-dimensional rotational KGE models. To address this, we exploit the orthogonality of rotation matrices, substituting rotation matrices($\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$) with orthogonal matrices ($\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$) of corresponding dimensions $d$. Consequently, our relation embedding ($\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$) are composed of $n/d$ block-diagonal orthogonal matrices $\mathbf{X}_{i}$ as illustrated in Figure 3:
@@
Where the number of relation parameters is $\frac{d(d-1)}{2}*\frac{n}{d}=\frac{(d-1)n}{2}$. And this aspect allows OrthogonalE to gain generality, adapting to datasets with diverse complexities by modifying the block-diagonal matrices’ dimension dd."
@@ 4.2 Matrices Representation for Entities
To enhance OrthogonalE’s flexibility, we aim to regulate entity dimension using variable $m$ and transform entity vectors $\mathbf{e}_{v}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ as shown in Figure 3, thus preventing unnecessary expansion of the relation size. This part allows OrthogonalE to acquire flexibility, adapting to diverse datasets with varying relation and entity parameters, rather than indiscriminately increase both. And the number of entity parameters is nmn*m.
@@ 4.3 Scoring function and Loss
4.4 Optimization
Initially, when optimizing relations, all entity parameters are fixed, rendering the entity embeddings analogous to the function $f(\cdot)$ in the problem 1. Notably, each orthogonal matrix $\mathbf{X}_{i}$ within the relation embedding $\mathbf{e}_{R}$ optimized by individual Riemannian optimization using RiemannianAdam Kochurov et al. (2020), which is a Riemannian version (equation 2) of the popular Adam optimizer Kingma and Ba (2014). These are then concatenated in a block-diagonal way according to equation 3 to complete the process. After optimizing the relation parameters $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$, they are held constant while the entity parameters $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ are optimized using Stochastic Gradient Descent (SGD), specifically employing the Adagrad optimizer Duchi et al. (2011).
5 Experiment
We expect that our proposed OrthogonalE model, employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations, will outperform baseline models. Also, we anticipate OrthogonalE is a general and flexible KGE model that saves parameters while effectively balancing high-dimensional block-diagonal orthogonal matrices $\mathbf{X}_{i}$ and entity matrices $\mathbf{e}_{V}$. Our goal is to validate these through empirical testing.
5.1 Experiment Setup
Dataset Entities Relations Train Validation Test
WN18RR 40,943 11 86,835 3,034 3,134
FB15k-237 14,541 237 272,115 17,535 20,466
Table 1: Details of the two datasets.
WN18RR FB15K-237
@@ 5.1 Experiment Setup
Gram-Schmidt($\times2$) .475 .434 .489 .556 .317 .226 .344 .502
Gram-Schmidt($\times3$) .487 .445 .500 .568 .322 .232 .350 .504
OrthogonalE($\times2$) .490 .445 .503 .573 .330 .239 .368 .516
OrthogonalE($\times3$) .493 .450 .508 .580 .331 .240 .359 .513
Table 2: Link prediction accuracy results of two datasets, Bold indicates the best score, and underline represents the second-best score. For a fair comparison, we standardized $m$ at 1 for Gram-Schmidt and all OrthogonalE sizes. The entity dimension for WN18RR was set to approximately 500 (e.g. 501 for $\times3$ blocks to ensure experimental feasibility), and around 1000 for FB15K-237. $\diamondsuit$]: Results are taken from Zhang et al. (2019). The result of RotatESun et al. (2019) and QuatEZhang et al. (2019) are reported without self-adversarial negative sampling and type-constrains respectively for fair comparison.

### v2

@@ Abstract
The primary aim of Knowledge Graph Embeddings (KGE) is to learn low-dimensional representations of entities and relations for predicting missing facts. Although rotation-based methods like RotatE Sun et al. (2019) and QuatE Zhang et al. (2019) perform well in KGE, they face two challenges: limited model flexibility requiring proportional increases in relation size with entity dimension, and difficulties in generalizing the model for higher-dimensional rotations. To address these issues, we introduce OrthogonalE, a novel KGE model employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. This approach not only enhances the generality and flexibility of KGE models but also captures several relation patterns that rotation-based methods can identify. Experimental results indicate that our new KGE model, OrthogonalE, offers generality and flexibility, captures several relation patterns, and significantly outperforms state-of-the-art KGE models while substantially reducing the number of relation parameters. Our code is available at OrthogonalE
@@ 1 Introduction
The fundamental elements of knowledge graphs (KGs) are factual triples, each represented as $(h,r,t)$, indicating a relationship $r$ between head entity $h$ and tail entity $t$. Notable examples include Freebase Bollacker et al. (2008), Yago Suchanek et al. (2007), and WordNet Miller (1995). KGs have practical applications in various fields such as question-answering Hao et al. (2017), information retrieval Xiong et al. (2017), recommender systems Zhang et al. (2016), and natural language processing Yang and Mitchell (2019), garnering considerable interest in academic and commercial research.
@@
Recently, rotation-based KGE methods have achieved notable success in the field. For instance, RotatE Sun et al. (2019) conceptualizes relations as 2D rotations while QuatE Zhang et al. (2019) employs 3D rotations to obtain a more expressive model than RotatE. Essentially, as illustrated in Fig. 1 , both operate by multiplying the relation matrix $\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}$ composed of Block-Diagonal Rotation matrix $\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$ (RotatE: $\mathbb{R}^{2\times 2}$, QuatE: $\mathbb{R}^{3\times 3}$) with the head entity vector $\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}$.
Figure 1: Fundamental operations ($\mathbf{e}_{R}^{1}\cdot\mathbf{e}_{h}^{1}\approx\mathbf{e}_{t}^{1}$) and inherent challenges of rotation-based KGE models. Rotation-based methods require increasing relation parameters for adequate entity representation and struggle with researching higher-dimensional rotation embeddings ($d>3$) due to their complexity. OrthogonalE, depicted in Fig. 2, efficiently resolves these challenges.
@@
Nevertheless, these approaches encounter two primary issues, as illustrated in Fig. 1. First, the model’s lack of flexibility necessitates increasing the overall relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{n\times n}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{(n+l)\times(n+l)}$) to meet entity dimension requirements ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{n}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{n+l}$) for better represent entities. For example, when the entity vector changes ($\mathbf{e}_{h}^{1}\in\mathbb{R}^{100}\rightarrow\mathbf{e}_{h}^{2}\in\mathbb{R}^{1000}$), the parameter increase is 900, but the corresponding change in the relation matrix ($\mathbf{e}_{R}^{1}\in\mathbb{R}^{100\times 100}\rightarrow\mathbf{e}_{R}^{2}\in\mathbb{R}^{1000\times 1000}$) results in a parameter increase of 990,000. This substantial increase leads to redundancy and inefficiency in representing relations.
Second, exploring high-dimensional rotational KGE models is challenging due to the significant computational demands and complexity of rotations in higher dimensions ($\mathbf{B}_{i}:\mathbb{R}^{2\times 2},\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{d\times d},d>3$), such as SO(4), SO(5), and SO(10). This restricts the development of more generalized and higher-dimensional rotation KGE approaches.
Figure 2: Diagram of the OrthogonalE approach. We employ matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations, thereby retaining the advantages of rotation-based method relation patterns while addressing its two main issues.
[…]
$\mathcal{V}$ Entity sets
$\mathcal{R}$ Relation sets
$\mathbf{e}_{v}\in\mathbb{R}^{n}$ Entity vector rep
$\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ Entity matrix rep in OrthogonalE
$\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ Relation matrix rep
$\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$ Block-Diagonal rotation matrix
$\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ Block-Diagonal orthogonal matrix
$n\in\mathbb{R}^{1}$ Row size of relation matrix rep
$m\in\mathbb{R}^{1}$ Column size of entity matrix rep
$d\in\mathbb{R}^{1}$ size of Block-Diagonal matrix
$d^{E}\left(.,.\right)$ Euclidean distance
$b_{v}\in\mathbb{R}^{1}$ Entity bias
$\cdot$ Matrix multiplication
$s(h,r,t)$ Scoring function
[…]
Table 1: Explanation of notations. Within the table, $\mathbf{e}_{v}$ includes the head $\mathbf{e}_{h}$ and tail $\mathbf{e}_{t}$ entity vectors as used in traditional KGE methods, whereas $\mathbf{e}_{V}$ consists of the head $\mathbf{e}_{H}$ and tail $\mathbf{e}_{T}$ entity matrix representations in our OrthogonalE approach. Furthermore, ’rep’ in the table denotes representation.
@@
To overcome these two issues, we propose a highly general and flexible KGE model named OrthogonalE as shown in Fig. 2, and detail of the notations has been shown in Table. 1. Firstly, by transforming entity vectors $\mathbf{e}_{v}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ for better represent entities, we control the entity dimension through variable $m$, avoiding unnecessary expansion of the relation size. Corresponding to the above example, we can maintain relation size ($\mathbf{e}_{R}\in\mathbb{R}^{100\times 100}$) and only modify entity matrix size ($\mathbf{e}_{H}^{1}\in\mathbb{R}^{100\times 1}\rightarrow\mathbf{e}_{H}^{2}\in\mathbb{R}^{100\times 10},m:1\rightarrow 10$) to meet the requirements of entity representation. Secondly, leveraging the concept that rotation matrices are orthogonal, we replace rotation matrices $\mathbf{B}_{i}$ with orthogonal matrices $\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$ of adaptable dimensions $d$, facilitating the exploration of higher-dimensional block-diagonal orthogonal matrix models. Lastly, for effective optimization, we employ Riemannian optimization for the relation matrix $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$ and Stochastic Gradient Descent (SGD) for the entity matrix $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$.
@@
We evaluate the new model on two KGE datasets including WN18RR Dettmers et al. (2018), FB15K-237 Toutanova and Chen (2015). Experimental results indicate that our new KGE model, OrthogonalE, offers generality and flexibility, captures several relation patterns, and significantly outperforms state-of-the-art KGE models while substantially reducing the number of relation parameters.
@@ 2 Related Work
Translation-based approaches are prominent in KGE, notably TransE Bordes et al. (2013), which interprets relations as vector translations. TransH Wang et al. (2014), TransR Lin et al. (2015), and TransD Ji et al. (2015) represent extensions of the translation-based method, building upon the foundational approach of TransE. ComplExTrouillon et al. (2016) advances this by embedding KGs in a complex space and using the Hermitian product for modeling antisymmetric patterns. Inspired by ComplEx, RotatE Sun et al. (2019) then innovated by treating relations as rotations in a complex vector space, capable of capturing varied relation patterns like Symmetry, Antisymmetry, Inversion, and Commutative Composition. Following this, QuatE Zhang et al. (2019) employed quaternion operations (3D rotations) for even better expressiveness than RotatE. DensE Lu et al. (2022) employed various techniques for 3D rotation implementation and proposed that 3D rotation could handle the relation pattern of non-commutative composition. HopfE Bastos et al. (2021) seeks to employ SO(4) rather than SO(3) for KG representation, which is directly connected to the generality issue discussed in our research. We are also keen on investigating rotations in higher dimensions. Nonetheless, progressing to SO(5) or even SO(10) poses substantial difficulties.
In conclusion, considering the two major disadvantages of rotation-based methods mentioned in the Introduction 1, there is a need to refine our model to make it more general and flexible.
@@
Meanwhile, we often use Riemannian Optimization for the orthogonal manifold, which has also progressed in deep learning, especially in CNNs and RNNs. Cho and Lee (2017) innovatively substituted CNN’s Batch Normalization layers with Riemannian optimization on the Grassmann manifold for parameter normalization. Additionally, significant strides in stabilizing RNN training have been made by Vorontsov et al. (2017), Wisdom et al. (2016), Lezcano-Casado and Martınez-Rubio (2019), and Helfrich et al. (2018), through the application of Riemannian optimization to unitary matrices.
@@ 3.2 Optimization on the orthogonal manifold
To address problem 1 more efficiently, recent studies suggest optimization on the orthogonal manifold with retractions as an effective approach Ablin and Peyré (2022). In this work, we primarily employ the retraction with exponential map method for iterative optimization, as illustrated in Fig. 3. The key iteration formula for this method is:
@@
Where $\text{expm}(\cdot)$ denotes the matrix exponential.
@@ 4 Approach
Our approach is developed to acquire both a flexible and general KGE model and ensure that this model can concurrently represent several relation patterns. This is achieved by employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations. Figure 2 illustrates the OrthogonalE approach, and Table 1 provides the details of the notations used.
@@ 4.1 Orthogonal Matrices for Relations
[…]
QuatE $n$ $n$ $n*m$
OrthogonalE($d$) $\frac{(d-1)n}{2}$ $\frac{(d-1)n}{2}$ $\frac{(d-1)n}{2}$
[…]
QuatE $n$ $n$ $n*m$
OrthogonalE($d$) $n*m$ $n$ $n*m$
[…]
Table 2: The parameter calculations for the KGE models. For all models, the relation matrix size is $n$. The block-diagonal matrix size is 2 for RotatE, 3 for QuatE, and $d$ for OrthogonalE($d\times d$), with an entity matrix column size of $m$ for OrthogonalE. In the table, "Normal" represents the standard parameter calculation, "$m=1$" constrains the column size of the entity matrix to 1 to explore the impact of block-diagonal orthogonal matrices on the model, as analyzed in section 5.2.2. "Uniform Entity" ensures that the entity dimensions are consistent across all models to demonstrate the parameter savings in the relation matrix when using the entity matrix in OrthogonalE, as discussed in section 5.2.4.
To address the challenge of exploring high-dimensional rotational KGE models mentioned in the introduction. We exploit the orthogonality of rotation matrices, substituting rotation matrices ($\mathbf{B}_{i}\in\mathbb{R}^{d\times d}$) with orthogonal matrices ($\mathbf{X}_{i}\in\mathbb{R}^{d\times d}$) of corresponding dimensions $d$. Consequently, our relation embedding ($\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$) are composed of $n/d$ block-diagonal orthogonal matrices $\mathbf{X}_{i}$ as illustrated in Fig. 2:
@@
Where the number of relation parameters is $\frac{d(d-1)}{2}*\frac{n}{d}=\frac{(d-1)n}{2}$, which shown in Table. 2. And this aspect allows OrthogonalE to gain generality, adapting to datasets with diverse complexities by modifying the block-diagonal matrices’ dimension $d$. Additionally, the employed relation structure facilitates the model’s capability to concurrently capture Symmetry, Antisymmetry, Inversion, and Non-commutative Composition relation patterns, as substantiated in Appendix A.2, and detailed introduction of relation patterns refer to Appendix A.4.
@@ 4.2 Matrices Representation for Entities
Inspired by Miyato et al. (2022), transforms vector embeddings into matrix embeddings to enhance embedding effectiveness. In our work, to enhance OrthogonalE’s flexibility, we aim to regulate entity dimension using variable $m$ and transform entity vectors $\mathbf{e}_{v}\in\mathbb{R}^{n}$ into matrices $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ as shown in Fig. 2, thus preventing unnecessary expansion of the relation size. This part allows OrthogonalE to acquire flexibility, adapting to diverse datasets with varying relation and entity parameters, rather than indiscriminately increase both. And the number of entity parameters is $n*m$.
@@ 4.3 Scoring function and Loss
$y_{t^{\prime}}=\left\{\begin{array}[]{l}-1,\text{ if }t^{\prime}=t\\ 1,\text{ otherwise }\end{array}\right.$ (7)
@@
HopfE Bastos et al. (2021) .472 .413 .500 .586 .343 .247 .379 .534
DensE Lu et al. (2022) .486 - - .572 .306 - - .481
[…]
Gram-Schmidt(3$\times$3) .487 .445 .500 .568 .322 .232 .350 .504
OrthogonalE(2$\times$2) .490 .445 .503 .573 .330 .239 .368 .516
OrthogonalE(3$\times$3) .493 .450 .508 .580 .331 .240 .359 .513
OrthogonalE(4$\times$4) .493 .446 .506 .578 .332 .240 .363 .517
OrthogonalE(10$\times$10) .494 .446 .508 .573 .334 .242 .367 .518
[…]
Table 3: Link prediction accuracy results of two datasets, Bold indicates the best score, and underline represents the second-best score. For a fair comparison, we standardized $m$ at 1 for Gram-Schmidt and all OrthogonalE sizes. The entity dimension for WN18RR was set to approximately 500 (e.g. 501 for 3$\times$3 blocks to ensure experimental feasibility), and around 1000 for FB15K-237. [$\diamondsuit$]: Results are sourced from Zhang et al. (2019). For a fair comparison, the results of RotatE, QuatE, HopfE, and DensE are reported without self-adversarial negative sampling, type constraints, semantics, or reciprocals. More baseline results have been shown in Appendix A.5.
4.4 Optimization
Initially, when optimizing relations, all entity parameters are fixed, rendering the entity embeddings analogous to the function $f(\cdot)$ in the problem 1. Notably, each block-diagonal orthogonal matrix $\mathbf{X}_{i}$ within the relation embedding $\mathbf{e}_{R}$ optimized by individual Riemannian optimization using RiemannianAdam Kochurov et al. (2020), which is a Riemannian version (equation 2) of the popular Adam optimizer Kingma and Ba (2014). These are then concatenated in a block-diagonal way according to equation 3 to complete the process. After optimizing the relation parameters $\mathbf{e}_{R}\in\mathbb{R}^{n\times n}$, they are held constant while the entity parameters $\mathbf{e}_{V}\in\mathbb{R}^{n\times m}$ are optimized using Stochastic Gradient Descent (SGD), specifically employing the Adagrad optimizer Duchi et al. (2011).
5 Experiment
We expect that our proposed OrthogonalE model, employing matrices for entities and block-diagonal orthogonal matrices with Riemannian optimization for relations, will outperform baseline models. Also, we anticipate OrthogonalE is a general and flexible KGE model and it can represent several relation patterns simultaneously. Our goal is to validate these through empirical testing.
5.1 Experiment Setup
[…]
FB15K-237 14,541 237 272,115 17,535 20,466
[…]
Table 4: Details of the two datasets.